Sol Summary Report – February 28th

Crew 275 Sol Summary Report 28Feb2023

Sol: 16

Summary Title: An extraordinary routine

Author’s name: Jeremy Rabineau, Commander.

Mission Status: Nominal.

Sol Activity Summary:

With a probability of precipitation of 0%, I must say that we were all surprised to see that the surroundings were covered with snow when we woke up. Since no EVA was planned for today, we took some time to relax a bit in the morning with a longer sport session including 30 minutes of yoga.

With the anchors of the location tracking system fully deployed, we have started equipping the crew with tags that they have to always keep on their arm or on their belt. This system enables us to check who is in which part of the station at which time. Later on, it will allow the scientists to correlate these locations and the environmental conditions at these locations with other parameters such as stress levels extracted from heart rate variability.

With the yesterday’s harvest, everyone is delighted to have the opportunity to eat fresh food. Our Crew Botanist and our Health & Safety Officer have prepared some delicious meals that recharged our batteries. We then discussed about more profound topics, which we did not really have the opportunity to do so far. These small things are extremely important for the mood, as we have already entered our third week here, and the end of the mission still looks relatively far.

In the afternoon, the crewmembers assigned to tomorrow’s EVA have taken some time to prepare for their mission with a 3D render of Candor Chasma. Tomorrow, they will only be allowed to read their notes to find the checkpoints defined during the previous EVA there.

Our Crew Astronomer also took the opportunity of today’s weather to visually observe the Sun. Unfortunately, he could not record any video of his observation. On his first attempts, the Sun hid behind some clouds, while on the second attempt, it hid behind the landscape, as it was already getting late.

Look Ahead Plan:

We plan an EVA at Candor Chasma for tomorrow (Sol 17) and another on Friday (Sol 19). The EVA crews will have to find specific checkpoints based on a 2D map or a 3D render. We are also planning to have a second geology EVA on Sol 20, to further test the MetMet instrument. We would like to take the opportunity of these different EVAs to change the batteries of the weather instruments, since they cannot last for more than two days, especially in cold weather.

The BBC is planned to arrive at the MDRS on Friday (Sol 19) between 8:30 and 9:00. They would follow us on our EVA to Candor Chasma, then film inside the station in the afternoon, and leave at around 17:00.

Anomalies in work: None.

Weather: Windy, low -5°C/23°F, high 7°C/44°F.

Crew Physical Status: Optimal.

EVA: None.

Reports to be filed: Sol Summary, EVA Request, Operations Report, Journalist

Report, GreenHab Report, Astronomy Report.

Support Requested: None.

Copyright © The Mars Society. All rights reserved. | Main Site